# Conditionally Site-Independent Neural Evolution of Antibody Sequences

### 06 Mar 2026 v2

@@ 3 Related Work
To move beyond the independent-sites assumption, a common approach consists of conditioning the parameterization of site specific mutation rates on the surrounding sequence context. This idea was first explored by Felsenstein1996AHM who used a Hidden Markov Model to assign per-site rate categories depending on adjacent sites. Brard2008SolvableMO proposed a neighbor-dependent DNA substitution model that adds extra terms to a baseline rate matrix in the presence of CpG sites. More recently, Albors2025APA used a neural network that takes a DNA sequence as input and outputs the parameters of an F81 model (Felsenstein2005EvolutionaryTF) for the central position of this sequence.
@@ 4 CoSiNE: Conditionally Site-Independent Neural Evolution Model
To address the limitations of classical models, we developed the CoSiNE model, which introduces two key modifications that mitigate model misspecification and leverage high-dimensional neural parameterization. First, we decouple rate estimation by learning site-specific rate matrices $Q_{\ell}$ for each site $\ell$ instead of a single unified matrix with scaling factors. Second, we model each evolutionary transition $\tau=(x,y,t)$ with its own set of rate matrices $\{Q^{(\tau)}_{\ell}\}_{\ell=1}^{L}$, which are inferred by conditioning on the parent sequence $x$. Concretely, our transition probability of CoSiNE is parameterized by:
@@
where $Q_{\theta}:\mathbb{R}^{|\mathcal{A}|^{L}}\to\mathbb{R}^{L\times|\mathcal{A}|\times|\mathcal{A}|}$ is a function parameterized by a neural network. Intuitively, CoSiNE relaxes the standard independent-sites assumption by estimating state-conditional rates, which enables the model to learn epistatic effects. In the next section, we provide theoretical grounding for CoSiNE by showing that CoSiNE constitutes a first-order approximation of a sequential point mutation process over the full sequence space. We further justify this framework by establishing an upper bound for the error between the transition probability of our model and the transition probability of the point mutation process $P(y\mid x,t)=\exp(t\mathbf{Q})_{xy}$, motivating its application to affinity maturation. Finally, we propose to sample from CoSiNE using a Gillespie procedure that provably samples from $P(y\mid x,t)$ under certain conditions.
@@
The validity of Section˜4 and Section˜4 relies on $(Q_{\theta}(x)_{\ell})_{x_{\ell},y_{\ell}}$ approximating the instantaneous non-zero rates in the point mutation process rate matrix $\mathbf{Q}_{x,y}$. However, because CoSiNE is trained with the approximate transition probability in Equation˜2, its estimated rates will incur some bias due to the marginalization of unobserved intermediate states on long branches. We acknowledge this theoretical limitation of CoSiNE and study its effect thoroughly in Section˜C.1. In practice, we show that the transition probability error of our model indeed scales quadratically with the branch length and that Gillespie sampling produces final state distributions that match those of the true process much more closely (Sections˜C.2 and C.3).
@@ 5.3 Zero-Shot Variant Effect Prediction
We investigate the efficacy of our approach to correcting for SHM on zero-shot variant effect prediction performance. As shown in Figure˜4, our selection score leads to increased correlation with fitness across all DMS datasets. Figure˜S6 in Section˜C.5 provides an intuitive picture of how the SHM correction helps isolate a selection signal from CoSiNE. We also see that the ability to model inter-chain interactions is important. In Section˜C.6 we ablate the VEP experiments by passing in just the heavy or light chain alone, finding that the selection score’s correlation with fitness drops substantially for some datasets.
@@
[…]
@@
ProGen2-Small 0.407 $\phantom{+}\underline{0.513}$ 0.024-0.024 $\phantom{+}0.098$ $\phantom{+}0.332$ 0.119 0.070
@@
[…]
CoSiNE per-site 0.630   0.520 $\phantom{+}\underline{0.453}$ $\phantom{+}\underline{0.442}$ $\phantom{+}\underline{0.361}$ 0.519 0.526

### 15 Mar 2026 v3

@@ 3 Related Work
To move beyond the independent-sites assumption, a common approach consists of conditioning the parameterization of site specific mutation rates on the surrounding sequence context. This idea was first explored by Felsenstein1996AHM who used a Hidden Markov Model to assign per-site rate categories depending on adjacent sites. Brard2008SolvableMO proposed a neighbor-dependent DNA substitution model that adds extra terms to a baseline rate matrix in the presence of CpG sites. More recently, Albors2025APA used a neural network that takes an entire DNA sequence as input and outputs the parameters of an F81 model (Felsenstein2005EvolutionaryTF) for the central position of this sequence.
@@ 4 CoSiNE: Conditionally Site-Independent Neural Evolution Model
To address the limitations of classical models, we developed the CoSiNE model, which introduces two key modifications that mitigate model misspecification and leverage high-dimensional neural parameterization. First, we decouple rate estimation by learning site-specific rate matrices $Q_{\ell}$ for each site $\ell$ instead of a single unified matrix with scaling factors. Second, we model each evolutionary transition $\tau=(x,y,t)$ with its own set of rate matrices $\{Q^{(\tau)}_{\ell}\}_{\ell=1}^{L}$, which are inferred by conditioning on the parent sequence $x$. Concretely, the transition probability of CoSiNE is given by
@@
where $Q_{\theta}:\mathbb{R}^{|\mathcal{A}|^{L}}\to\mathbb{R}^{L\times|\mathcal{A}|\times|\mathcal{A}|}$ is a function parameterized by a neural network. Intuitively, CoSiNE relaxes the standard independent-sites assumption by estimating state-conditional rates, which enables the model to learn epistatic effects. In the next section, we provide theoretical grounding for CoSiNE by showing that our model constitutes a first-order approximation of a sequential point mutation process over the full sequence space. We further justify this framework by establishing an upper bound for the error between the transition probability of our model and the transition probability of the point mutation process $P(y\mid x,t)=\exp(t\mathbf{Q})_{xy}$. Finally, we propose to sample from CoSiNE using a Gillespie procedure that provably samples from $P(y\mid x,t)$ under certain conditions.
@@
The validity of Section˜4 and Section˜4 relies on $(Q_{\theta}(x)_{\ell})_{x_{\ell},y_{\ell}}$ approximating the instantaneous non-zero rates in the point mutation process rate matrix $\mathbf{Q}_{x,y}$. However, because CoSiNE is trained with the approximate transition probability in Equation˜2, its estimated rates will incur some bias due to the marginalization of unobserved intermediate states on long branches. We acknowledge this theoretical limitation and study its effect thoroughly in Section˜C.1. In practice, we show that the transition probability error of our model indeed scales quadratically with the branch length and that Gillespie sampling produces final state distributions that match those of the true process much more closely (Sections˜C.2 and C.3).
@@ 5.3 Zero-Shot Variant Effect Prediction
We investigate the efficacy of our approach to correcting for SHM on zero-shot VEP performance. As shown in Figure˜4, our selection score leads to increased correlation with fitness across all datasets. Figure˜S6 provides an intuitive picture of how the SHM correction helps isolate a selection signal from CoSiNE. We also see that the ability to model inter-chain interactions is important. In Section˜C.6, we ablate this ability by passing in just the heavy or light chain alone, finding that the selection score’s correlation with fitness drops substantially for some datasets.
@@
DASM 0.596 $\phantom{+}0.474$ $\phantom{+}\underline{0.270}$ $\phantom{+}\underline{0.415}$ $\phantom{+}0.327$ 0.450 0.536
@@
ProGen2-Small 0.407 $\phantom{+}\mathbf{0.513}$ 0.024-0.024 $\phantom{+}0.098$ $\phantom{+}\underline{0.332}$ 0.119 0.070
@@
CoSiNE (ours) 0.613 $\phantom{+}\underline{0.508}$ $\phantom{+}\mathbf{0.464}$ $\phantom{+}\mathbf{0.456}$   0.371 0.498 0.536